OpenStudy (anonymous):

Subtract the following radicals: 12√3 - 7√27

jigglypuff314 (jigglypuff314):

radicals (or square roots) are like variables in that you can only combine those that have common bases, like in 2x + 3y + 2x = 4x + 3y but you can't add the 4x and the 3y for your problem, you can look at it this way: 12√3 - 7√27 = 12*√3 - 7*√(9*3) = 12*√3 - 7*√9 *√3 = 12√3 - 7*3*√3 = 12√3 - 21√3 then you would get - 9√3